Ms. McGuillicuddy is painting her house fern green. She needs 15 gallons of paint. The formula for mixing fern green is 4 parts yellow to 1 part blue . How many gallons of the yellow paint and of the blue paint will Ms. McGuillicuddy need?
Answer by rothauserc(4718)

let x be gallons of paint, then
4x + x = 15
5x = 15
x = 3
using the original equation, we know Ms. McGuillicuddy needs
4 * 3 = 12 gallons of yellow paint
1 * 3 = 3 gallons of blue paint